OpenStudy (anonymous):

What is the solution to the rational equation (x/2x-1) - (1/4) = (2/2x-1)? a) x=2 b) x= 7/2 c) x=3 d) x=7/6 I think its c but Im not sure c)x=

OpenStudy (anonymous):

take lcm on left hand side and solve for x

OpenStudy (anonymous):

how do i do that?

OpenStudy (anonymous):

yu can confirm whether your answer is right or wrong by substituting the value of x on both sides of the equarion if LHS=RHS,the answer is true

OpenStudy (anonymous):

((4x)-(2x-1))/(4(2x-1))=2/(2x-1) (2x+1)/(4(2x-1))=2/(2x-1) now cancelling the common factor in denominator on both sides,we get 2x+1/4=2 2x+1=8 2x=7 x=7/2

OpenStudy (anonymous):

thank you! I worked it out and its true :)

OpenStudy (anonymous):

you r welcome............so you re my fan!!!!!!!!!!!!!